SX On-Sky Checkout

07:17
  • Getting Started.
  • catalog_show is not showing the image of the guidestars.
07:45
  • Presetting on AO32 = HD1480
  • Preset failed at optimize gain, switching to Intervention Mode.
  • Failed when I resent the preset, then the pointing messed up.
  • Jenny says that the failure I saw in my first preset was possibly from the system starting optical gain before the pupils were well centered. Whatever the case, it applied TT in one direction and then went to "Apply Optical Gain", which of course failed.
8:00
  • Preset successful. Starting NCPA.
  • Before starting NCPA, I'll take some datasets without.
  • 080221, 080246, 080313, 080355, 080422
  • Cube W/LUCI 10 frames, no NCPA 082234
  • Cube W NCPA 082551
  • Missed the Dark.
  • Going to run LUCI Script
  • NCPA ON!
8:48
  • AO91 - LHS 1443
  • Datasets:
  • Oops, it's bin3
  • Shell ripped! 8:55 Unknown cause
09:04
  • timeout setting flat
  • Fast diagnostic and adsecarb were down
09:07
  • AO115
  • Forcing Bin4. It wants to run in bin 3, so I set it to bin4, 250Hz and skipped Check flux. Didn't work.
  • None of the AO stars are faint enough.
  • Jenny is looking for one.
9:33
  • Star 0 of Stone A1 is 17mag
09:37:56
  • Actuator jumped the instant that we sent the preset. I think that it was due to an Ice exception, a series of them were reported by AOS.
09:38
  • Too much lunar background
  • Looking for a different one
  • Stone A1-17
  • Re-homed the rerotator it was spinning out of control.
  • Found the wrong star, it was 15th mag instead of 17th
  • Still looking
10:00
  • Failure-user error, tried to do steps out of order on the flao gui, unknown as to why it failed after safe-skipping.
  • I think that we have too much background, we're too close to the moon and the WFS is behaving erratically. (I get 16.3 mag when we are pointed at empty sky)
  • I checked the cameras and we are pointed almost directly at it.
10:32
  • 17.47 mag star
  • Guider shows 18.1
  • Camera lens has hit its limit.
  • Having Josh rotate the LUCI rotator +90
10:43
  • We are on the star, camera lens is centering...slowly
  • We are showing 18.0 mag
  • 3.8 photons per subap
  • 105009, 105106
  • Testing Offsets, n' stuff
  • Images are kind of Elongated, but 18.0!
10:44
  • 13.7 star in M52
  • Bin2 849 Hz
  • GOPT is On
  • Dataset: 111533, 111621, 111640
  • Loop opened, GOPT often makes bin2 unstable.
  • Resending preset
-- GregoryTaylor - 04 Sep 2020
Topic revision: r3 - 04 Sep 2020, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback